OpenStudy (anonymous):

Solve the system. y = -1/3 x + 2 and x + 3y = 3

OpenStudy (anonymous):

have you tried the problem yourself?

OpenStudy (anonymous):

\(\frac{1}{3}x - y=-2\) multi by -3 \[x+3y=3\] -x+3y=6 +x+3y=3 add two equation 6y = 9 y = 3/2 sub y in first equation -x+(3/2)3=6 -x+9/2=6 x=-3/2